Determine the molecular formula of a substance that has the following proximate formula: 40.9% carbon, 4.5% hydrogen, and 54.6% oxygen. Data: H-1, C-12, O-16.


1
Expert's answer
2021-06-21T05:31:43-0400

Solution:

Suppose the sample has a mass of 100 g.

Convert the %values to grams:

Mass of C = w(C) × Mass of sample = 0.409 × 100 g = 40.9 g C

Mass of H = w(H) × Mass of sample = 0.045 × 100 g = 4.5 g H

Mass of O = w(O) × Mass of sample = 0.546 × 100 g = 54.6 g O


Convert grams to moles:

40.9 g C × (1 mol C / 12 g C) = 3.41 mol C

4.5 g H × (1 mol H / 1 g H) = 4.50 mol H

54.6 g O × (1 mol O / 16 g O) = 3.41 mol O


Divide all moles by the smallest of the results:

C: 3.41 / 3.41 = 1.00

H: 4.50 / 3.41 = 1.32

O: 3.41 / 3.41 = 1.00


This means that: C1H1.3O1

Since we are looking for the smallest whole number ratio that exists between the elements, multiply each subscript by 3 to get the empirical formula:

3 × (C1H1.3O1) ⇒ C3H4O3

Therefore, the empirical formula is C3H4O3

To determine the molecular formula, you need to know the molar mass of of a substance:

Molar mass / Empirical formula mass = n

Thus, the general view of the molecular formula is (C3H4O3)n

In the case of n = 1, the molecular formula will be C3H4O3

In the case of n = 2, the molecular formula will be (C3H4O3)2 ⇒ C6H8O6